OpenStudy (anonymous):

A triangle with side lengths 33 feet, 56 feet and 65 feet is a right triangle. true or false

OpenStudy (texaschic101):

in a right triangle A^2 + B^2 = C^2...and C is the largest number so lets check.. A = 33 , B = 56, and C = 65 A^2 + B^2 = C^2 33^2 + 56^2 = 65^2 work the rest out..and if both sides come out equal, then it is a right triangle